# A Communication Consistent Approach to Signal Temporal Logic Task Decomposition in Multi-Agent Systems

### v2

@@ 6.1.2 Cost Function
Let a single inconsistent task $\phi_{ij}=\varphi_{ij}$ with truth set $\mathcal{B}_{ij}$ be decomposed over the path $\bm{\pi}_{i}^{j}$. By Lemma 1, the inclusion relation (32) must hold for a valid decomposition. By Prop. 5 we also know that (32) implies the relation (35) $\sum_{(r,s)\in\epsilon(\bm{\pi}_{i}^{j})}\alpha_{rs}^{ij}\leq 1$ (where the notation in place before the definition of the index mapping $y_{\Pi_{rs}}$ is used here and replaced in (49)), which holds with equality iff (32) does. Intuitively, the optimal decomposition is obtained when (32) holds with equality, as this is equivalent to recovering the truth set $\mathcal{B}_{ij}$ in its full extent, after the decomposition (cf. Prop. 5). Since this occurs only if $\sum_{(r,s)\in\epsilon(\bm{\pi}_{i}^{j})}\alpha_{rs}^{ij}=1$, then the decomposition accuracy in (35) is considered an appropriate metric to define ”optimality” for our proposed decomposition. The objective function of our decomposition becomes
@@
$\underset{y_{\Pi_{rs}}}{\equiv}\sum_{(r,s)\in\mathcal{E}_{\pi}}\sum_{l\in\bar{\mathcal{K}}_{rs}}\alpha_{rs}^{l},$
where the first equality is obtained by swapping an iteration over the inconsistent edges $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$, with an iteration over the decomposition edges $(r,s)\in\mathcal{E}_{\pi}$. On the other hand, the last equivalence is only notational though $y_{\Pi_{rs}}$ as per (38).
6.1.3 Constraints
The constraints for our optimization program are of three main types: communication consistency, conflicting conjunctions and decomposition constraints. Consider first the communication constraints over an edge $(r,s)\in\mathcal{E}_{\pi}$. For each parametric task $\varphi_{rs}^{l},\;l\in\bar{\mathcal{K}}_{rs}$ there is a parametric truth set $\mathcal{B}_{rs}^{l}=\mathcal{P}(A_{rs}^{l},\bm{c}_{rs}^{l},\alpha_{rs}^{l}\bm{z}_{rs}^{l})$ with generators set $V_{rs}^{l}=\nu(\mathcal{P}(A_{rs}^{l},\bm{0},\bm{z}_{rs}^{l}))$. Then, by Prop. 5.10, communication consistency is enforced by (47). We can repeat the same argument for each $l\in\bar{\mathcal{K}}_{rs}$ to obtain:
@@ 6.1.3 Constraints
where the matrices $N_{k}^{l},\;\forall k=1,\ldots|V_{rs}^{l}|$ are as per (48). The second type of constraint guarantees conflict-free conjunctions over the tasks $\phi_{rs}$ as per (40). In this respect, consider, for each edge $(r,s)\in\mathcal{E}_{\pi}$, the set $\mathfrak{Q}_{rs}$ as per (44). The set $\mathfrak{Q}_{rs}$ is the set of subsets of indices in $\mathcal{K}_{rs}$ such that, for each $Q\in\mathfrak{Q}_{rs}$, a conjunction of tasks $\land_{l\in Q}\varphi_{rs}^{l}$ over $(r,s)\in\mathcal{E}_{\pi}$ could cause a conflict. Then, conflicting conjunctions are avoided by imposing (46) over each edge $(r,s)\in\mathcal{E}_{\pi}$. The third, and last, type of constraint is the one required to obtain a valid decomposition of each inconsistent task $\phi_{ij}=\varphi_{ij},\;\forall(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ by the satisfaction of (33). This is achieved by the satisfaction of (33) and (34). Differently from the previous two types of constraints, this constraint is shared among multiple edges $(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$ in the sense that its satisfaction involves parameters vectors spanning all the edges $(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$ along the path $\bm{\pi}_{i}^{j}$. With the intent of uniforming the notation via the index mapping $y_{\Pi_{rs}}$, we let $l_{rs}=y_{\Pi_{rs}}((i,j))$ for each $(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$ to be the unique index associated with the inconsistent edge $(i,j)$ over the edge $(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$, such that the constraint (33) can be rewritten as
@@
6.1.4 Optimization program
By letting $\mathcal{H}=\bigcup_{(r,s)\in\epsilon(\bm{\omega})}\mathcal{H}_{rs}$ and $\Xi=\bigcup_{(r,s)\in\epsilon(\bm{\omega})}\Xi_{rs}$ the centralised optimization program applied to solve the task decomposition problem is given by:
@@ 6.1.4 Optimization program
Remark 6.12.
The proposed task decomposition approach is agnostic to the chosen dynamics in (1). Indeed, the decomposition derived as the solution to (50) depends only on the geometry of the predicate level sets of the tasks and their logical relation (e.g., potential conflicting conjunctions). Future work will address the interaction between dynamics, communication, and STL constraints to enhance the generality of the proposed decomposition framework.
6.2 A decentralized solution
@@
where the functions $f_{rs}(\bm{\chi}_{rs})=\sum_{l\in\bar{\mathcal{K}}_{rs}}\alpha_{rs}^{l},\;\forall(r,s)\in\mathcal{E}_{\pi},\;\forall(r,s)\in\mathcal{E}_{\pi}$ are convex and the sets $\mathcal{C}_{rs},\;\forall(r,s)\in\mathcal{E}_{\pi}$ are compact and convex sets. Moreover, $T_{rs}\in\mathbb{R}^{(\sum_{(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}m_{ij})\times((n+1)\cdot|\mathcal{H}_{rs}|+n\cdot|\mathfrak{Q}_{rs}|)}$ and $\bm{t}_{rs}\in\mathbb{R}^{(\sum_{(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}m_{ij})}$ with $m_{ij}$ being the number of rows of each matrices $M_{ij},\forall(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ as per (34).
Proof 6.14.
@@
First note that $f_{rs}(\bm{\chi}_{rs}),\forall(r,s)\in\mathcal{E}_{\pi}$, are linear (and thus convex) functions since they are obtained as sums of elements of the vectors $\bm{\chi}_{rs}$. Next, we prove 1) the equivalence between (50a) and (52a), 2) the equivalence between (50b)-(50c)-(50e) and (52b), 3) the equivalence between (50d) and (52c). The equivalence 1) of the cost function (52a) and (50a) is evident since $f_{rs}(\bm{\chi}_{rs})=\sum_{l\in\mathcal{K}_{rs}}\alpha_{rs}^{l}$. The equivalence 2) follows by considering, for each $(r,s)\in\mathcal{E}_{\pi}$, the sets $\mathcal{C}^{1}_{rs}=\{\bm{\chi}_{rs}\,|\,(\bm{\eta}_{rs}^{l})^{T}N^{l}_{k}(\bm{\eta}_{rs}^{l})\leq r^{2}_{c},\forall l\in\bar{\mathcal{K}}_{rs}\;k=1,\ldots|V_{rs}^{l}|\}$, $\mathcal{C}^{2}_{rs}=\{\bm{\chi}_{rs}\,|\,A_{rs}^{l}\bm{\xi}_{rs}^{q}+\left[A_{rs}^{l}\mid\bm{z}_{rs}^{l}\vphantom{X}\right]\bm{\eta}_{rs}^{l}\leq\bm{0},\;\forall l\in\bar{\mathcal{K}}_{rs},\forall Q\in\mathfrak{Q}_{rs},\;\text{s.t},\;q=y_{\mathfrak{Q}_{rs}(Q)}\}$, $\mathcal{C}^{3}_{rs}=\{\bm{\chi}_{rs}\,|\,\|\bm{\xi}_{rs}^{q}\|\leq\xi_{max}\;\forall q=1,\ldots|\mathfrak{Q}_{rs}|\land\|\bm{\eta}_{rs}^{l}\|\leq\eta_{max}\;\forall l\in\bar{\mathcal{K}}_{rs}\},$ which are, by definition, equivalent to (50b)-(50c)-(50e) when considered for a single edge $(r,s)\in\mathcal{E}_{\pi}$. Since $\mathcal{C}^{1}_{rs}$, $\mathcal{C}^{2}_{rs}$ and $\mathcal{C}^{3}_{rs}$ are convex, and $\mathcal{C}^{3}_{rs}$ is bounded, then $\mathcal{C}_{rs}=\mathcal{C}^{1}_{rs}\cap\mathcal{C}^{2}_{rs}\cap\mathcal{C}^{3}_{rs}$ is also convex and bounded, thus proving the equivalence. The last equivalence between (50d) and (52c) is proved as follows. Consider, for each $(r,s)\in\mathcal{E}_{\pi}$, the matrix $W_{rs}\in\mathbb{R}^{(\sum_{(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}m_{ij})\times(n+1)\cdot|\Pi_{rs}|}$, where we recall that $\Pi_{rs}$ contains all the inconsistent edges $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ whose decomposition path is passing by $(r,s)\in\mathcal{E}_{\pi}$ as per (36b) and, by definition, $|\bar{\mathcal{K}}_{rs}|=|\Pi_{rs}|=|\mathcal{H}_{rs}|$. Moreover $m_{ij}$ is the number of rows of the matrix $M_{ij}$ associated with each inconsistent each edge $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$. We define the matrices $W_{rs}$ by blocks as $W_{rs}[p,w]$ with row-blocks $p=1,\ldots|\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}|$ and $w=1,\ldots|\Pi_{rs}|$ column-blocks such that all the matrices $W_{rs}$ have the same number of row-blocks, but only a number of column-blocks equal to the number of edges $|\Pi_{rs}|$ that are decomposed over the edge $(r,s)\in\mathcal{E}_{\pi}$. We define the one-to-one index mapping $y_{\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}:\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}\rightarrow\{1,\ldots|\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}|\}$ associating each inconsistent edges $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ with a unique row-block index. Then each matrix $W_{rs}$ has non-zero blocks only at those $(p,w)$ index pairs such that $(i,j)=y_{\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}^{-1}(p)$ is an inconsistent edge and the mapping $y_{\Pi_{rs}}^{-1}(w+K_{rs})=(i,j)$, as per (38), is well defined, meaning that the decomposition path $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ passes though $(r,s)$. Thus, $W_{rs}$ is defined by blocks as
@@
With these definitions, if we let $\bm{o}_{rs}=T_{rs}\bm{\chi}_{rs}-\bm{t}_{rs}$, then $\bm{o}_{rs}$ is a vector defined by $|\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}|$ row-blocks of dimension $\mathbb{R}^{m_{ij}}$ such that
@@
where we recall that $l_{rs}=y_{\Pi_{rs}}((i,j))$ is the index assigned to the inconsistent edge $(i,j)$ over the edge $(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$ applied for the decomposition, which is different from the index $l_{r^{\prime}s^{\prime}}=y_{\Pi_{r^{\prime}s^{\prime}}}((i,j))$ assigned to the same edge $(i,j)$ by another edge $(r^{\prime},s^{\prime})\in\epsilon(\bm{\pi}_{i}^{j})$ applied for the decomposition. Eventually, noting that $\sum_{(r,s)\in\mathcal{E}_{\pi}}\bm{o}_{rs}$ is equivalent to (50d) concludes the proof.
Given the optimization problem (50), a solution can be found in a decentralized fashion by leveraging the fact that (50) has the edge-wise decentralized structure in (52). We propose to employ the algorithm in [29] to solve (52) in a decentralized fashion as explained next. First, we wish to define an edge-computing graph denoted as $\mathcal{G}_{\Theta}$ such that each node in $\mathcal{G}_{\Theta}$ corresponds to an edge $(r,s)\in\mathcal{E}_{\pi}$. Hence, the set of nodes for $\mathcal{G}_{\Theta}$ is given by $\mathcal{V}_{\theta}=\mathcal{E}_{\pi}$. On the other hand, the set of edges for $\mathcal{G}_{\Theta}$ is given by
$\mathcal{E}_{\Theta}=\{((r,s),(r^{\prime},s^{\prime}))\in\mathcal{V}_{\Theta}\times\mathcal{V}_{\Theta}\;|\;\{r,s\}\cap\{r^{\prime},s^{\prime}\}\neq\emptyset\}.$ (54)
In other words, a tuple $((r,s),(r^{\prime},s^{\prime}))$ is an edge in $\mathcal{E}_{\Theta}$ if $(r,s)\in\mathcal{E}_{\pi}$ shares a node with $(r^{\prime},s^{\prime})\in\mathcal{E}_{\pi}$. Then the edge-computing graph is defined as $\mathcal{G}_{\Theta}(V_{\Theta},\mathcal{E}_{\Theta})$ and the neighbour set is defined as $\mathcal{N}_{\Theta}((r,s))=\{(r^{\prime},s^{\prime})\in\mathcal{V}_{\Theta}|((r,s),(r^{\prime},s^{\prime}))\in\mathcal{E}_{\Theta}\}$. The edge-computing graph $\mathcal{G}_{\Theta}$ can then be applied to find a solution for (52) in a decentralized fashion where each node $(r,s)\in\mathcal{V}_{\Theta}$ (an edge in $\mathcal{E}_{\pi}$) solves iteratively the following local problem iteratively

### v3

@@ VI-A2 Cost Function
Let a single inconsistent collaborative task $\phi_{ij}=\varphi_{ij}$ with truth set $\mathcal{B}_{ij}$ be decomposed over the path $\bm{\pi}_{i}^{j}$. By Lemma 1, the inclusion relation (33) must hold for a valid decomposition, and by Prop. 5 we already argued that the satisfaction of (33) implies the relation (36) $\sum_{(r,s)\in\epsilon(\bm{\pi}_{i}^{j})}\alpha_{rs}^{ij}\leq 1$ (where the notation in place before the definition of the index mapping $y_{\Pi_{rs}}$ is used here and replaced in (50)), which holds with equality if and only if (33) does. Intuitively, the optimal decomposition is indeed obtained when (33) holds with equality as this is equivalent to recovering the truth set $\mathcal{B}_{ij}$ in its full extent, after the decomposition. Since this condition occurs only when $\sum_{(r,s)\in\epsilon(\bm{\pi}_{i}^{j})}\alpha_{rs}^{ij}=1$, then the decomposition accuracy in (36) is considered further as an appropriate metric to define ”optimality” for our proposed decomposition. Therefore, the objective function of our decomposition becomes
@@
$=\sum_{(r,s)\in\mathcal{E}_{\pi}}\sum_{l\in\bar{\mathcal{K}}_{rs}}\alpha_{rs}^{l},$
where the first equality is obtained by swapping an iteration over the inconsistent edges $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$, with an iteration over the decomposition edges $(r,s)\in\mathcal{E}_{\pi}$. On the other hand, the last equality is only notational and obtained by leveraging the index map $y_{\Pi_{rs}}:\Pi_{rs}\mapsto\bar{\mathcal{K}}_{rs}$ as per (39).
VI-A3 Constraints
The constraints for our optimization program are of three main types: communication consistency, conflicting conjunctions and decomposition constraints. Consider first the communication constraints over an edge $(r,s)\in\mathcal{E}_{\pi}$. For each parametric task $\varphi_{rs}^{l},\;l\in\bar{\mathcal{K}}_{rs}$ there is a parametric truth set $\mathcal{B}_{rs}^{l}=\mathcal{P}(A_{rs}^{l},\bm{c}_{rs}^{l},\alpha_{rs}^{l}\bm{z}_{rs}^{l})$ with generators set $V_{rs}^{l}=\nu(\mathcal{P}(A_{rs}^{l},\bm{0},\bm{z}_{rs}^{l}))$. Then, by Prop. 9, communication consistency is enforced by (48). We can repeat the same argument for each $l\in\bar{\mathcal{K}}_{rs}$ to obtain :
@@ VI-A3 Constraints
where the matrices $N_{k}^{l},\;\forall k=1,\ldots|V_{rs}^{l}|$ are as per (49).
The second type of constraint guarantees conflict-free conjunctions over the tasks $\phi_{rs}$ as per (41). In this respect, consider, for each edge $(r,s)\in\mathcal{E}_{\pi}$, the set $\mathfrak{Q}_{rs}$ as per (45). The set $\mathfrak{Q}_{rs}$ is the set of subsets of indices in $\mathcal{K}_{rs}$ such that, for each $Q\in\mathfrak{Q}_{rs}$, a conjunction of tasks $\land_{l\in Q}\varphi_{rs}^{l}$ over $(r,s)\in\mathcal{E}_{\pi}$ could cause a conflict. Then, conflicting conjunctions are avoided by imposing (47) over each edge $(r,s)\in\mathcal{E}_{\pi}$.
The third, and last, type of constraint is the one required to obtain a valid decomposition of each inconsistent task $\phi_{ij}=\varphi_{ij},\;\forall(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ by the satisfaction of (34). This is achieved by the satisfaction of (34) and (35). Differently from the previous two types of constraints, this constraint is shared among multiple edges $(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$ in the sense that its satisfaction involves parameters vectors spanning all the edges $(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$ along the path $\bm{\pi}_{i}^{j}$. With the intent of uniforming the notation via the index mapping $y_{\Pi_{rs}}$, we let $l_{rs}=y_{\Pi_{rs}}((i,j))$ for each $(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$ to be the unique index associated with the inconsistent edge $(i,j)$ over the edge $(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$, such that the constraint (34) can be rewritten as
@@
VI-A4 Optimization program
By letting $\mathcal{H}=\bigcup_{(r,s)\in\epsilon(\bm{\omega})}\mathcal{H}_{rs}$ and $\Xi=\bigcup_{(r,s)\in\epsilon(\bm{\omega})}\Xi_{rs}$ the centralised optimization program applied to solve the task decomposition problem is then given by:
@@ VI-A4 Optimization program
VI-B A decentralized solution
@@
where the functions [^3] The functions $f_{rs}$ should not be confused with the system dynamics as it is clear from the context $f_{rs}(\bm{\chi}_{rs})=\sum_{l\in\mathcal{K}_{rs}}\alpha_{rs}^{l},\;\forall(r,s)\in\mathcal{E}_{\pi},\;\forall(r,s)\in\mathcal{E}_{\pi}$ are convex and the sets $\mathcal{C}_{rs},\;\forall(r,s)\in\mathcal{E}_{\pi}$ are compact and convex sets. Moreover, $T_{rs}\in\mathbb{R}^{(\sum_{(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}m_{ij})\times((n+1)\cdot|\mathcal{H}_{rs}|+n\cdot|\Xi_{rs}|)}$ and $\bm{t}_{rs}\in\mathbb{R}^{(\sum_{(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}m_{ij})}$ with $m_{ij}$ is the number of rows of each matrix $M_{ij}$ as per (35).
Proof:
@@
First note that $f_{rs}(\bm{\chi}_{rs}),\forall(r,s)\in\mathcal{E}_{\Pi}$, are linear (and thus convex) functions since they are obtained as sums of elements of the vectors $\bm{\chi}_{rs}$. Next, we prove 1) the equivalence between (51a) and (53a), 2) the equivalence between (51b)-(51c)-(51e) and (53b), 3) the equivalence between (51d) and (53c). The equivalence 1) of the cost function (53a) and (51a) is evident since $f_{rs}(\bm{\chi}_{rs})=\sum_{l\in\mathcal{K}_{rs}}\alpha_{rs}^{l}$. The equivalence 2) follows by considering, for each $(r,s)\in\mathcal{E}_{rs}$, the sets $\mathcal{C}^{1}_{rs}=\{\bm{\chi}_{rs}\,|\,(\bm{\eta}_{rs}^{l})^{T}N^{l}_{k}(\bm{\eta}_{rs}^{l})\leq r^{2}_{c},\forall l\in\bar{\mathcal{K}}_{rs}\;k=1,\ldots|V_{rs}^{l}|\}$, $\mathcal{C}^{2}_{rs}=\{\bm{\chi}_{rs}\,|\,A_{rs}^{l}\bm{\xi}_{rs}^{q}+\left[A_{rs}^{l}\mid\bm{z}_{rs}^{l}\vphantom{X}\right]\bm{\eta}_{rs}^{l}\leq\bm{0},\;\forall l\in\bar{\mathcal{K}}_{rs},\forall Q\in\mathfrak{Q}_{rs},\;\text{s.t},\;q=y_{\mathfrak{Q}_{rs}(Q)}\}$, $\mathcal{C}^{3}_{rs}=\{\bm{\chi}_{rs}\,|\,\|\bm{\xi}_{rs}^{q}\|\leq\xi_{max}\;\forall q=1,\ldots|\mathfrak{Q}_{rs}|\land\|\bm{\eta}_{rs}^{l}\|\leq\eta_{max}\;\forall l\in\bar{\mathcal{K}}_{rs}\},$ which are equivalent to (51b)-(51c)-(51e) when considered for a single edge $(r,s)\in\mathcal{E}_{\pi}$. Since $\mathcal{C}^{1}_{rs}$, $\mathcal{C}^{2}_{rs}$ and $\mathcal{C}^{3}_{rs}$ are convex, and $\mathcal{C}^{3}_{rs}$ is bounded, then $\mathcal{C}_{rs}=\mathcal{C}^{1}_{rs}\cap\mathcal{C}^{2}_{rs}\cap\mathcal{C}^{3}_{rs}$ is also convex and bounded, thus proving the equivalence. The last equivalence between (51d) and (53c) is proved as follows. Consider, for each $(r,s)\in\mathcal{E}_{\pi}$, the matrix $W_{rs}\in\mathbb{R}^{(\sum_{(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}m_{ij})\times(n+1)\cdot|\Pi_{rs}|}$, where we recall that $\Pi_{rs}$ contains all the inconsistent edges $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ whose decomposition path is passing by $(r,s)\in\mathcal{E}_{\pi}$ as per (37b) and $|\Pi_{rs}|=|\bar{\mathcal{K}}_{rs}|$ by (38). Moreover $m_{ij}$ is the number of rows of the matrix $M_{ij}$ associated with each inconsistent each edge $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$. We define the matrices $W_{rs}$ by blocks as $W_{rs}[p,w]$ with row-blocks $p=1,\ldots|\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}|$ and $w=1,\ldots|\Pi_{rs}|$ column-blocks such that all the matrices $W_{rs}$ have the same number of row-blocks, but only a number of column-blocks equal to the number of edges $|\Pi_{rs}|$ that are decomposed over the edge $(r,s)\in\mathcal{E}_{\pi}$ (and recall that $|\bar{\mathcal{K}}_{rs}|=|\Pi_{rs}|$). We define the one-to-one index mapping $y_{\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}:\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}\rightarrow\{1,\ldots|\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}|\}$ associating each inconsistent edges $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ with a unique row-block index. Then each matrix $W_{rs}$ has non-zero blocks only at those $(p,w)$ index pairs such that $(i,j)=y_{\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}^{-1}(p)$ is an inconsistent edge and the mapping $y_{\Pi_{rs}}^{-1}(w+K_{rs})=(i,j)$, as per (39), is well defined, meaning that the decomposition path $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ passes though $(r,s)$. Thus, $W_{rs}$ is defined by blocks as
@@
With these definitions, if we let $\bm{o}_{rs}=T_{rs}\bm{\chi}_{rs}-\bm{t}_{rs}$, then $\bm{o}_{rs}$ is a vector defined by a number of $|\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}|$ blocks of dimension $\mathbb{R}^{m_{ij}}$ such that
@@
where the notation $l_{rs}$ is introduced again to be consistent with (51d) and make clear that $l_{rs}=y_{\Pi_{rs}}((i,j))$ is the index assigned to the inconsistent edge $(i,j)$ over $(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$, which is different from the index $l_{r^{\prime}s^{\prime}}=y_{\Pi_{r^{\prime}s^{\prime}}}((i,j))$ assigned to the same edge $(i,j)$ by another edge $(r^{\prime},s^{\prime})\in\epsilon(\bm{\pi}_{i}^{j})$. Eventually, noting that $\sum_{(r,s)\in\mathcal{E}_{\pi}}\bm{o}_{rs}$ is equivalent to (51d) concludes the proof. ∎
Given the optimization problem (51), a solution can be found in a decentralized fashion by leveraging the fact that (51) has the edge-wise decentralized structure in (53). We propose to employ the algorithm in [29] to solve (53) in a decentralized fashion as explained next. First, we wish to define an edge-computing graph denoted as $\mathcal{G}_{\Theta}$ such that each node in $\mathcal{G}_{\Theta}$ corresponds to an edge $(r,s)\in\mathcal{E}_{\pi}$. Hence, the set of nodes for $\mathcal{G}_{\Theta}$ is given by $\mathcal{V}_{\theta}=\mathcal{E}_{\pi}$. On the other hand, the set of edges for $\mathcal{G}_{\Theta}$ is given by $\mathcal{E}_{\Theta}=\{((r,s),(r^{\prime},s^{\prime}))\in\mathcal{V}_{\Theta}\times\mathcal{V}_{\Theta}\;|\;\{r,s\}\cap\{r^{\prime},s^{\prime}\}\neq\emptyset\}$. In other words, a tuple $((r,s),(r^{\prime},s^{\prime}))$ is an edge in $\mathcal{E}_{\Theta}$ if $(r,s)\in\mathcal{E}_{\pi}$ shares a node with $(r^{\prime},s^{\prime})\in\mathcal{E}_{\pi}$. Then the edge-computing graph is defined as $\mathcal{G}_{\Theta}(V_{\Theta},\mathcal{E}_{\Theta})$ and the neighbour set is defined as $\mathcal{N}_{\Theta}((r,s))=\{(r^{\prime},s^{\prime})\in\mathcal{V}_{\Theta}|((r,s),(r^{\prime},s^{\prime}))\in\mathcal{E}_{\Theta}\}$. The edge-computing graph $\mathcal{G}_{\Theta}$ can then be applied to find a solution for (53) in a decentralized fashion where each node $(r,s)\in\mathcal{V}_{\Theta}$ (an edge in $\mathcal{E}_{\pi}$) solves the following relaxed local problem iteratively